# Helical $f$-Wave Superconductivity in Cubic Rashba Superconductors

### v2

@@ Abstract
The Rashba effect–a momentum-dependent spin splitting from broken inversion symmetry–is a crucial component for engineering topological superconductors (TSCs). However, time-reversal invariant TSCs based on conventional linear Rashba spin-orbit coupling (RSOC) are fundamentally restricted to a helical $p$-wave state, hosting only a single Majorana Kramers pair. In this Letter, we demonstrate that the unique triple-winding spin texture of cubic RSOC offers a direct pathway to surpass this limitation. We propose a bilayer model where the synergy of cubic RSOC and an intrinsic odd-parity pairing instability yields a rare 2D helical $f$-wave TSC. This state is characterized by a large mirror Chern number (MCN) ${\cal N}_{\text{M}}\!\!=\!\!3$ and hosts three Kramers pairs of Majorana edge modes. Furthermore, the interplay of linear and cubic RSOCs can yield a hybrid helical $p+f$-wave TSC with an even larger MCN of ${\cal N}_{\text{M}}\!\!=\!\!4$. Our work establishes higher-order spin-orbit coupling as a powerful design principle for realizing TSCs with multiple Majorana Kramers channels, holding immediate relevance for tunable platforms like oxide heterostructures.
@@
Spin-orbit coupling is the relativistic interaction between an electron’s spin and its orbital motion. In crystals lacking inversion symmetry, this interaction manifests as Rashba spin-orbit coupling (RSOC), which lifts the spin degeneracy of electronic bands through momentum-dependent spin splitting [1]. RSOC is the cornerstone for realizing exotic quantum phenomena [2, 3, 4, 5, 6]. In superconductors, RSOC facilitates the admixture of spin singlet and triplet Cooper pairs [7], significantly enhancing the upper critical field [7, 8, 9]. Crucially, RSOC extends beyond noncentrosymmetric materials. A hidden RSOC can emerge even in globally centrosymmetric structures, provided that local atomic environments break inversion symmetry [10, 11, 12]. This phenomenon, observed in systems such as bismuth-based cuprates  [11] and heavy fermion superconductors [12], has significantly broadened the materials landscape and intensified the search for novel superconducting phases, particularly those characterized by odd-parity Cooper pairing [13].
@@
A primary motivation for studying RSOC is its vital role in engineering topological superconductors (TSCs)—exotic states of matter that host Majorana quasiparticles. Seminal proposals have leveraged the spin-momentum locking of conventional, linear-in-momentum ($\bm{k}$) RSOC to generate TSCs. For instance, an $s$-wave superconductor proximitized to the surface of a topological insulator can realize an effective $p_{x}+ip_{y}$ state supporting Majorana zero modes in vortex cores [14]. By breaking time-reversal symmetry (TRS) with a magnetic field, two-dimensional (2D) systems with linear RSOC can host chiral $p$-wave superconductivity [15, 16, 17, 18]. Furthermore, linear RSOC is central to proposals for 2D helical TSCs preserving TRS. These states, classified in symmetry class DIII by a $\mathbb{Z}_{2}$ topological invariant [19, 20, 21, 22], support a protected Kramers pair of Majorana modes counterpropagating along each edge. Crucially, this $\mathbb{Z}_{2}$ classification fundamentally restricts such systems to only a single Kramers pair of Majorana modes. A crucial question, therefore, is how to transcend this binary framework to realize TSCs with multiple Kramers pairs of Majorana channels, effectively transitioning from a $\mathbb{Z}_{2}$ to a $\mathbb{Z}$ classification.
Figure 1: (a, b) Spin texture on the spin-split FSs, illustrating (a) a single spin winding for linear RSOC and (b) a triple winding for cubic RSOC. (c, d) Helical Majorana channels in a Rashba bilayer connected by mirror symmetry, under dominant (c) linear RSOC and (d) cubic RSOC.
@@
A promising avenue involve exploring spin textures beyond the simple winding of the linear RSOC. Recently, significant attention has shifted towards a higher-order RSOC [23, 24, 25, 26, 27, 28, 29, 30, 25, 31, 32, 33, 34], cubic in $\bm{k}$, which can dominate when the linear term is suppressed by symmetry. This cubic RSOC is particularly relevant in oxide interfaces such as LaAlO3/SrTiO3 (LAO/STO), where its strength is electrically tunable [28, 35] and superconductivity is well established [36, 37]. The resulting spin textures are markedly different. Unlike linear RSOC, where the spin winds once around the Fermi surface (FS) [Fig. 1(a)], cubic RSOC generates a complex, triple-winding spin texture: as an electron encircles the FS, its spin rotates three times [Fig. 1(b)]. While prior work has explored proximity-induced, TRS-breaking TSCs [38, 39, 40], the potential for a TRS-invariant TSC driven by an intrinsic pairing instability in a cubic RSOC system–and the consequent realization of multiple Majorana Kramers pairs arising from its unique triple-winding spin texture–remain key open questions.
@@
In this Letter, we provide an affirmative answer and establish a direct mechanism for realizing a novel TRS invariant helical $f$-wave TSC that hosts multiple Kramers pairs of Majorana channels. We study an interacting bilayer model with local inversion symmetry breaking–a canonical platform for hidden Rashba physics. We find that the triple-winding spin texture inherent to cubic RSOC, combined with an odd-parity pairing instability, leads to a TSC protected by mirror symmetry. The bulk topology is characterized by a large mirror Chern number (MCN), ${\cal N}_{\text{M}}\!\!=\!\!3$, guaranteeing three Kramers pairs of Majorana edge modes [Fig. 1(d)]. This contrasts sharply with the helical $p$-wave state found in mirror-symmetric bilayers with only linear RSOC [20, 22], which host only a single pair of such modes [Fig. 1(c)]. Furthermore, we show that the interplay of linear and cubic RSOCs yields a hybrid helical p+fp\!\!+\!\!f TSC with ${\cal N}_{\text{M}}\!\!=\!\!4$.
@@
${\cal H}_{\pm i}(\bm{k})=\begin{pmatrix}{\cal H}_{\text{N},\pm i}(\bm{k})+\mu&-i\Delta_{0}s_{y}\\ i\Delta_{0}s_{y}&-{\cal H}^{\ast}_{\text{N},\pm i}(-\bm{k})-\mu\end{pmatrix}$.
@@
$\Delta_{++}(\bm{k})=-\Delta_{0}\frac{\alpha_{\text{CR}}}{\lambda_{\bm{k}}}\Big{(}-ik_{x}+k_{y}\Big{)}^{3},$
@@
This result is central to our work: the triple-winding spin texture of the normal state, encoded in the cubic momentum dependence of $\bm{g}_{\text{CR}}(\bm{k})$, is directly inherited by the intraband pairing potential. The resulting gap functions, $\Delta_{++}(\bm{k})\propto(k_{y}-ik_{x})^{3}$ and $\Delta_{--}(\bm{k})\propto(k_{y}+ik_{x})^{3}$ correspond to $f_{y(3x^{2}-y^{2})}\mp if_{x(x^{2}-3y^{2})}$ symmetries with opposite chiralities for the upper and lower bands. When $\mu$ lies within the HG [Fig. 2(b)], only the lower band crosses $\mu$. The system is then effectively a single band superconductor. In this regime, the ${\cal H}_{+i}$ sector describes a chiral $f$-wave TSC. The ${\cal H}_{-i}$ sector describes the time-reversed partner with opposite chirality. Together, they form a TRS-invariant TSC with helical $f$-wave pairing protected by mirror symmetry ${\cal M}_{z}$.
@@
Figure 3: FSs with spin texture, phase diagram and spectral function ${\cal A}(E,\bm{k})$ of the $\Delta_{3}$ pairing phase in a semi-infinite geometry. (a) FSs with spin texture for $\mu_{B}=-0.2$ and $\mu_{C}=-0.4$, top panel: pure cubic RSOC $\alpha_{\text{CR}}=0.1$, $\alpha_{\text{LR}}=0$. Bottom panel: coexisting linear and cubic RSOCs $\alpha_{\text{CR}}=0.1,\alpha_{\text{LR}}=0.1$. (b) ${\cal A}(E,k_{x})$ for cubic RSOC at $\mu_{B}=-0.2$, showing three pairs of helical Majorana modes. (c) Phase diagram of coexiting linear and cubic RSOCs at $\mu_{C}=-0.4$. (d) ${\cal A}(E,k_{x})$ for coexisting linear and cubic RSOCs at $\mu_{C}=-0.4$, showing four pairs of helical Majorana modes. Common parameters: $\beta_{0}=0.5,\beta_{1}=-1,\varepsilon=0.35,\Delta_{0}=0.2$.
@@
Since each mirror subsector belongs to class D, its topology is characterized by an integer Chern number, ${\cal N}_{\pm i}$. The overall topology of the system is classified by the MCN ${\cal N}_{\text{M}}\!\!=\!\!({\cal N}_{+i}\!-\!{\cal N}_{-i})/2$ [44, 45, 46, 47, 48, 49, 50]. A direct calculation shows that when $\mu$ is inside the HG and the system is fully gapped (requiring $\varepsilon^{2}>\mu^{2}+\Delta_{0}^{2}$), the MCN is ${\cal N}_{\text{M}}\!\!=\!\!3$. The magnitude ${\cal N}_{\text{M}}$ directly reflects the triple-winding topology of the cubic Rashba spin texture [Top panel of Fig. 3(a)], which has been faithfully transferred to the superconducting order parameter. This demonstrates the “topology multiplier” principle: the threefold winding of the spin texture generates a threefold enhancement of the topological invariant. The bulk-boundary correspondence dictates that a TSC with MCN ${\cal N}_{\text{M}}$ will host $|{\cal N}_{\text{M}}|$ pairs of counterpropagating Majorana edge modes. To verify this, we discretize the model on a square lattice and calculate the spectral function ${\cal A}(E,\bm{k})$ for a semi-infinite geometry. As shown in Fig. 3(b), ${\cal A}(E,\bm{k})$ for the $\Delta_{3}$ phase with ${\cal N}_{\text{M}}\!\!=\!\!3$ clearly shows three pairs of linearly dispersing, gapless modes crossing zero energy. These are the predicted three helical Majorana channels, in stark contrast to the single pair expected from conventional linear RSOC [19, 20, 21].
@@
For $\mu_{C}$, which features two concentric FSs enclosing the $\Gamma$ point, $\Delta_{3}$ remains favored when $U>0$ and $V<0$. Applying the established topological criteria for odd-parity superconductors [51, 41], the topological invariant is found to be trivial [43]. This is a direct consequence of the even number of FSs enclosing the TRS invariant momentum. This scenario is also observed in pure linear RSOC bilayer [20]. However, this classification is fundamentally altered when cubic and linear RSOCs coexist.
@@
Hybrid helical $p+f$-wave TSC.—A critical test for experimental relevance is the stability of a topological phase against perturbations expected in real materials. For instance, in oxide heterostructures like LAO/STO [28, 35], both linear and cubic RSOCs coexist, with their relative strengths tunable via carrier filling. We therefore investigated the robustness of the ${\cal N}_{\text{M}}\!\!=\!\!3$ phase by reintroducing the linear RSOC. We found that the three pairs of helical Majorana edge modes persist for dominant cubic RSOC as long as $\mu$ lies within HG (See SM[43]).
@@
Crucially, a new phenomenon emerges when $\mu$ is moved out of HG but remains within the lower band. This creates two distinct FSs: a smaller, inner surface near the $\Gamma$ point and a larger, outer surface farther away [Bottom panel of Fig. 3(a)]. In this regime $\Delta_{3}$ pairing remains favored under similar interaction conditions [Fig. 3(c)]. The interplay between coexisting linear and cubic RSOCs and $\Delta_{3}$ pairing induces an even richer TSC characterized by a heightened MCN, ${\cal N}_{\text{M}}\!\!=\!\!4$, hosting four Kramers pairs of Majorana edge modes [Fig. 3(d)]. This state is a beautiful example of $\bm{k}$-space topological engineering. The different momentum scaling of the two RSOCs allows them to dominate in different regions of the Brillouin zone. Given sufficient separation of the FSs in $\bm{k}$}-space, the linear term ($\propto k$) dominates the inner surface, fostering a helical $p$-wave state. Conversely, the cubic term ($\propto k^{3}$) dominates the outer surface, fostering a helical $f$-wave state. The resulting superconductor effectively merges these two distinct topologies, summing their respective contributions to the total MCN.

### v3

# Helical Superconductors with Multiple Majorana Kramers Pairs in Rashba Bilayers

@@ Abstract
The momentum dependence of Rashba spin-orbit coupling (RSOC) is a key ingredient for engineering topological superconductors (TSCs), but research has largely focused on the linear-in-momentum form. This focus has restricted time-reversal invariant TSCs to helical $p$-wave states with only a single Majorana Kramers pair, whose existence is tied to the criterion of an odd number of Fermi surfaces (FSs). In this Letter, we demonstrate that a bilayer system with pure cubic RSOC and intrinsic odd-parity pairing on a single FS yields a rare 2D helical $f$-wave TSC. This state is characterized by a large mirror Chern number (MCN) of ${\cal N}_{\text{M}}=3$ and hosts three Kramers pairs of Majorana edge modes. Remarkably, the interplay of linear and cubic RSOCs in this bilayer can generate a helical hybrid $p+f$-wave TSC with an even larger MCN of ${\cal N}_{\text{M}}=4$ from a normal state with two FSs, thereby circumventing the conventional odd-FS criterion. Our work establishes higher-order RSOC as a powerful design principle for realizing TSCs with multiple Majorana Kramers channels, fundamentally reshapes the helical hybrid for helical TSCs, and holds immediate relevance for tunable platforms like oxide heterostructures.
@@
Spin-orbit coupling is the relativistic interaction between an electron’s spin and its orbital motion. In crystals lacking inversion symmetry, this interaction manifests as Rashba spin-orbit coupling (RSOC), which lifts the spin degeneracy of electronic bands through momentum-dependent spin splitting [1]. RSOC is a cornerstone for realizing exotic quantum phenomena [2, 3, 4, 5, 6]. In superconductors, RSOC facilitates the admixture of spin-singlet and triplet Cooper pairs [7], significantly enhancing the upper critical field [7, 8, 9]. Crucially, the influence of RSOC is not limited to globally noncentrosymmetric materials; a hidden RSOC can emerge even in centrosymmetric structures, provided that local atomic environments break inversion symmetry [10, 11, 12]. This phenomenon, observed in systems such as bismuth-based cuprates [11] and heavy fermion superconductors [12], has significantly broadened the materials landscape for novel superconducting phases, particularly those characterized by odd-parity Cooper pairing [13].
@@
A primary motivation for studying RSOC is its vital role in engineering topological superconductors (TSCs)—exotic states of matter that host Majorana quasiparticles. Seminal proposals have leveraged the spin-momentum locking of conventional, linear-in-momentum ($\bm{k}$) RSOC to generate TSCs. For instance, an $s$-wave superconductor proximitized to the surface of a topological insulator can realize an effective $p_{x}+ip_{y}$ state supporting Majorana zero modes in vortex cores [14]. By breaking time-reversal symmetry (TRS) with a magnetic field, two-dimensional (2D) systems with linear RSOC can host chiral $p$-wave TSCs [15, 16, 17, 18]. Furthermore, linear RSOC is central to proposals for 2D helical TSCs that preserve TRS, whose existence is tied to the criterion of an odd number of Fermi surfaces (FSs) [19, 20, 21]. These states belong to symmetry class DIII and are characterized by a $\mathbb{Z}_{2}$ topological invariant, which protects a single Kramers pair of counterpropagating Majorana edge modes [22, 23, 24, 25]. A crucial question, therefore, is how to transcend this binary framework to realize 2D TSCs with multiple Kramers pairs of Majorana channels, effectively transitioning from a $\mathbb{Z}_{2}$ to a $\mathbb{Z}$ classification.
Figure 1: Spin textures and their resulting helical Majorana channels. (a) Triple spin winding on a single FS due to cubic RSOC. (b) A combination of triple (outer FS) and single (inner FS) spin windings arising from the interplay of cubic and linear RSOC. (c, d) Helical Majorana edge modes in a mirror-symmetric bilayer induced by (c) cubic RSOC, resulting in three channels, and (d) coexisting linear and cubic RSOCs, which yields four channels.
@@
A promising avenue involves exploring spin textures beyond the simple winding of the linear RSOC. Recently, significant attention has shifted towards a higher-order, cubic-in-$\bm{k}$ RSOC [26, 27, 28, 29, 30, 31, 32, 33, 28, 34, 35, 36, 37], which can dominate when the linear term is suppressed by symmetry. This cubic RSOC is particularly relevant in oxide interfaces such as LaAlO3/SrTiO3 (LAO/STO), where its strength is electrically tunable [31, 38] and superconductivity is well established [39, 40]. The resulting spin textures are markedly different. Unlike linear RSOC, where the spin winds once around the FS [see spin texture of the inner FS in Fig. 1(b)], cubic RSOC generates a complex, triple-winding spin texture: as an electron encircles the FS, its spin rotates three times [Fig. 1(a)]. While prior work has explored proximity-induced, TRS-breaking TSCs [41, 42, 43], the potential for a TRS-invariant TSC driven by an intrinsic pairing instability in a cubic RSOC system–and the consequent realization of multiple Majorana Kramers pairs arising from its unique triple-winding spin texture–remain key open questions.
@@
In this Letter, we provide an affirmative answer and establish a direct mechanism for realizing novel TRS invariant helical TSCs that host multiple Kramers pairs of Majorana channels. We study an interacting bilayer model with local inversion symmetry breaking–a canonical platform for hidden RSOC. We find that the triple-winding spin texture inherent to cubic RSOC, combined with an odd-parity pairing instability on a single FS, leads to a TSC protected by mirror symmetry. The bulk topology is characterized by a large mirror Chern number (MCN), ${\cal N}_{\text{M}}\!=\!3$, guaranteeing three Kramers pairs of Majorana edge modes [Fig. 1(c)]. Furthermore, we show that the coexistence of linear and cubic RSOCs can generate a helical hybrid $p+f$-wave TSC characterized by ${\cal N}_{\text{M}}\!=\!4$, hosting four Kramers pairs of Majorana edge modes [Fig. 1(d)]. Crucially, this state emerges from a normal state with two FSs, thereby circumventing the established odd-FS criterion for helical TSCs  [19, 20, 21].
@@
${\cal H}_{\pm i}(\bm{k})=\begin{pmatrix}{\cal H}_{\text{N},\pm i}(\bm{k})+\mu&-i\Delta_{0}s_{y}\\ i\Delta_{0}s_{y}&-{\cal H}^{\ast}_{\text{N},\pm i}(-\bm{k})-\mu\end{pmatrix}$,
@@
$\Delta_{++}(\bm{k})=-\Delta_{0}\frac{\alpha_{\text{CR}}}{\lambda_{\bm{k}}}\Big(-ik_{x}+k_{y}\Big)^{3},$ (7)
@@
This result is central to our work: the triple-winding spin texture of the normal state, encoded in the cubic momentum dependence of $\bm{g}_{\text{CR}}(\bm{k})$, is directly inherited by the intraband pairing potential. The resulting gap functions, $\Delta_{++}(\bm{k})\propto(k_{y}-ik_{x})^{3}$ and $\Delta_{--}(\bm{k})\propto(k_{y}+ik_{x})^{3}$ correspond to $f_{y(3x^{2}-y^{2})}\mp if_{x(x^{2}-3y^{2})}$ symmetries with opposite chiralities for the upper and lower bands. When $\mu$ lies within the HG [Fig. 2(a)], only the lower band crosses $\mu$. The system is then effectively a single band superconductor. In this regime, the ${\cal H}_{+i}$ sector describes a chiral $f$-wave TSC. The ${\cal H}_{-i}$ sector describes the time-reversed partner with opposite chirality. Together, they form a TRS-invariant TSC with helical $f$-wave pairing protected by mirror symmetry ${\cal M}_{z}$.
@@
Figure 3: Superconducting phase diagram and spectral function ${\cal A}(E,\bm{k})$ for the $\Delta_{3}$ pairing phase in a semi-infinite geometry. (a, b) ${\cal A}(E,k_{x})$ for the pure cubic RSOC case ($\alpha_{\text{CR}}=0.1$). Three Kramers pairs of Majorana edge modes appear when the chemical potential is in HG as in (a) $\mu_{B}=-0.2$, but not outside it as in (b) $\mu_{C}=-0.6$. (c) Phase diagram of coexiting linear and cubic RSOCs $\alpha_{\text{CR}}=\alpha_{\text{LR}}=0.1$ at $\mu_{C}=-0.4$. (d) ${\cal A}(E,k_{x})$ at the point marked by red star in (c), revealing four pairs of helical Majorana modes. Common parameters: $\beta_{0}=0.5,\beta_{1}=-1,\varepsilon=0.35,\Delta_{0}=0.2$.
@@
Since each mirror subsector belongs to class D, its topology is characterized by an integer Chern number, ${\cal N}_{\pm i}$. The overall topology of the system is classified by the MCN ${\cal N}_{\text{M}}\!\!=\!\!({\cal N}_{+i}\!-\!{\cal N}_{-i})/2$ [46, 47, 48, 49, 50, 51, 52]. A direct calculation shows that when $\mu$ is inside the HG and the system is fully gapped (requiring $\varepsilon^{2}>\mu^{2}+\Delta_{0}^{2}$), the MCN is ${\cal N}_{\text{M}}\!\!=\!\!3$. The magnitude ${\cal N}_{\text{M}}$ directly reflects the triple-winding topology of the cubic Rashba spin texture [Fig. 1(a)], which has been faithfully transferred to the superconducting order parameter. This demonstrates the “topology multiplier” principle: the threefold winding of the spin texture generates a threefold enhancement of the topological invariant. The bulk-boundary correspondence dictates that a TSC with MCN ${\cal N}_{\text{M}}$ will host $|{\cal N}_{\text{M}}|$ pairs of counterpropagating Majorana edge modes. To verify this, we discretize the model on a square lattice and calculate the spectral function ${\cal A}(E,\bm{k})$ for a semi-infinite geometry. As shown in Fig. 3(a), ${\cal A}(E,\bm{k})$ for the $\Delta_{3}$ phase with ${\cal N}_{\text{M}}\!\!=\!\!3$ clearly shows three pairs of linearly dispersing, gapless modes crossing zero energy. These are the predicted three helical Majorana channels, in stark contrast to the single pair expected from conventional linear RSOC [22, 23, 24].
@@
For $\mu_{C}$, which features two concentric FSs enclosing the $\Gamma$ point, $\Delta_{3}$ remains favored when $U>0$ and $V<0$. Applying the established topological criterion for odd-parity superconductors [20, 19], this superconducting state is found to be topologically trivial [Fig. 3(b)]. This is a direct consequence of the even number of FSs enclosing the TRS invariant momentum. This scenario is also observed in pure linear RSOC bilayer [23]. However, this classification is fundamentally altered when cubic and linear RSOCs coexist.
@@
Helical hybrid $p+f$-wave TSC.—A critical test for experimental relevance is the stability of a topological phase against perturbations expected in real materials. For instance, in oxide heterostructures like LAO/STO [31, 38], both linear and cubic RSOCs coexist, with their relative strengths tunable via carrier filling. We therefore investigated the robustness of the ${\cal N}_{\text{M}}\!\!=\!\!3$ phase by reintroducing the linear RSOC. We found that the three pairs of helical Majorana edge modes persist for dominant cubic RSOC as long as $\mu$ lies within HG (See SM[45]).
@@
Crucially, a new phenomenon emerges when $\mu$ is moved out of HG but remains within the lower band. This creates two distinct FSs: a smaller, inner surface near the $\Gamma$ point and a larger, outer surface farther away [Fig. 1(b)]. In this regime $\Delta_{3}$ pairing remains favored under similar interaction conditions [Fig. 3(c)]. The interplay between coexisting linear and cubic RSOCs and $\Delta_{3}$ pairing induces an even richer TSC characterized by a heightened MCN, ${\cal N}_{\text{M}}\!\!=\!\!4$, hosting four Kramers pairs of Majorana edge modes [Fig. 3(d)]. This state is a beautiful example of $\bm{k}$-space topological engineering. The different momentum scaling of the two RSOCs allows them to dominate in different regions of the Brillouin zone. Given sufficient separation of the FSs in $\bm{k}$}-space, the linear term ($\propto k$) dominates the inner surface, fostering a helical $p$-wave state. Conversely, the cubic term ($\propto k^{3}$) dominates the outer surface, fostering a helical $f$-wave state. The resulting superconductor effectively merges these two distinct topologies, summing their respective contributions to the total MCN.